In a local bar, a customer slides an empty beer mug down the counter for a refill. The height of the counter is 1.30 m. The mug slides off the counter and strikes the floor 1.20 m from the base of the counter.

(a) With what velocity did the mug leave the counter?
m/s

(b) What was the direction of the mug's velocity just before it hit the floor?
Answer #1

a) d/t=v, h=1/2(9.8)t^2, t=sqrt(h/-4.9)
v=d/sqrt(h/-4.9)

v=1.2/sqrt (1.3/-4.9)

v=-2.35 m/s


b. Vx=d/sqrt(h/4.9), Vy=-9.8t = -9.8sqrt(h/4.9)
direction=
arctan((-9.8sqrt(1.3/4.9)) / (1.2/sqrt(1.3/4.9))) =-2.12